# Learning to Evolve: Bayesian-Guided Continual Knowledge Graph Embedding

### v2

@@ Abstract.
Knowledge Graph, Clustering, Continual Learning
copyright: acmlicenseddoi: XXXXXXX.XXXXXXXcopyright: noneccs: Computing methodologies Artificial intelligenceccs: Information systems Data mining
@@ 1. Introduction
Knowledge graphs (KGs)(Liang et al., 2024b)(Ji et al., 2021) have become indispensable tools for representing structured information, supporting a wide range of knowledge driven applications such as LLM based question answering(Pan et al., 2024)(Xu et al., 2025) and recommender systems(Zhao et al., 2024)(Yuan et al., 2025b). Knowledge graph embedding (KGE) aims to encode entities and relations in a KG as continuous vectors, and the quality of these embeddings directly affects downstream performance. Classic KGE models (Bordes et al., 2013)(Sun et al., 2019) perform excellently on static KGs, yet real world graphs usually evolve continuously. For example, the YAGO knowledge base evolved from YAGO3(Mahdisoltani et al., 2013) to YAGO4.5(Suchanek et al., 2024), with the number of entities soaring from about 4.6 million to 49 million and facts (triples) from 40 million to 132 million. For such continually evolving knowledge graphs (CKGs), existing KGE models must retrain the entire graph after each update—a computationally prohibitive and impractical strategy for large scale graphs like YAGO. Hence, this work focuses on CKGE models capable of learning CKGs efficiently.
@@
2. Related Work
2.1. Knowledge Graph Embedding
Early KGE models mainly follow the translation paradigm, such as TransE(Bordes et al., 2013) and TransH(Wang et al., 2014) in the real vector space, and their extensions to complex/quaternion spaces, including RotatE(Sun et al., 2019), QuatE(Zhang et al., 2019), HAKE(Zhang et al., 2020), SATKGC(Ko et al., 2025), MRME(Li et al., 2025a), SymAgent(Liu et al., 2025), HySAE(Li et al., 2025c), and GIE(Cao et al., 2022). These approaches introduced the idea of representing a triple $(h,r,t)$ through simple algebraic operations in a vector space. Since then, both the efficiency and the expressive power of KGE models have improved along multiple dimensions. RecPiece (Liang et al., 2024a) clusters relation centres to select anchors, reducing memory consumption while improving link prediction accuracy; RAAKGC(Yuan et al., 2025a) pulls queries of a pretrained language model toward relation aware neighbour anchors to construct more discriminative embeddings; FedE(Zhou et al., 2024) pioneers a poisoning attack on federated KG embedding, using inferred relations to craft stealthy updates that induce specific errors without degrading main task performance. Recently, KGE research has begun to leverage diffusion models to recast the task as conditional entity generation. For example, FDM(Long et al., 2024) trains a diffusion model that outputs the distribution of tail entities conditioned on the head entity and relation, thereby capturing multimodal relational semantics without relying on complicated probabilistic mechanisms. S2DN(Ma et al., 2025) proposes a semantic-aware denoising network and filters out unreliable interactions to enhance the structural reliability of target links.
2.2. Continual Knowledge Graph Embedding
As knowledge graphs (KGs) evolve, embedding models must learn new facts while avoiding catastrophic forgetting of previously acquired knowledge. Research on continual knowledge graph embedding (CKGE) has progressed notably in recent years. IncDE (Liu et al., 2024a) orders new triples hierarchically and employs layer-wise incremental distillation to preserve old knowledge; FastKGE (Liu et al., 2024b) accelerates updates with an adaptive low-rank adapter (IncLoRA), substantially shortening training time while isolating parameters for new snapshots; LKGE (Cui et al., 2023) combines a masked autoencoder with transfer regularization to extend embeddings across multiple snapshots; ETT-CKGE (Zhu et al., 2025) replaces costly node scoring with lightweight task-driven labels and aligns KG snapshots through simple matrix operations; CLKGE (Cao, 2024) couples continual learning with an energy-based manifold so that new and old knowledge reinforce each other; CMKGE (Song et al., 2024) introduces a biologically inspired dual inhibition–excitation mask to balance plasticity and stability; FMR (Zhu et al., 2024) rotates the parameter space based on Fisher information and assigns flexible regularization to further mitigate forgetting; ERPP (Yang et al., 2025) transfers relational paths across snapshots to shift from catastrophic forgetting to knowledge accumulation; and SAGE (Li et al., 2025b) adaptively expands embedding dimensions and uses dynamic distillation for CKGE.
@@ 3.3. Continual learning and Bayesian neural networks
Exact inference is usually intractable for deep networks, so an independent Gaussian (mean field) approximation is commonly adopted (Lee et al., 2024b; Ahn et al., 2019): $q_{t}(\boldsymbol{\omega})=\prod_{i=1}^{P}\mathcal{N}\!\bigl(\omega_{i};\mu_{t,i},\sigma_{t,i}^{2}\bigr)$, with $\theta_{t}=(\boldsymbol{\mu}_{t},\boldsymbol{\sigma}_{t})$ taken as the learnable variational parameters. The independence assumption greatly simplifies the KL divergence computation and the reparameterization gradient estimator, enabling iterative posterior updates. In practice, the reparameterization trick $\omega_{i}=\mu_{t,i}+\sigma_{t,i}\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,1)$ is often employed to estimate gradients.
@@ 4.1. Framework Overview
The overall framework of BAKE is shown in Figure 2. Its core idea is to formalize the continual knowledge graph embedding (CKGE) problem as a serialized Bayesian inference process. Given a knowledge graph sequence $\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{T}$ that evolves over time, the model receives a set of newly added triples $\Delta\mathcal{T}_{t}$ at each time snapshot $t$. BAKE works together through the following three modules: First, Bayesian-guided knowledge evolution uses the parameter posterior distribution of the previous snapshot t1t-1 as a priori, integrates new knowledge and quantifies uncertainty through Bayesian updating; second, the continual clustering module introduces fairness regularization to maintain the semantic consistency of the embedding space; finally, the joint optimization objective integrates the above modules to achieve a balance between new knowledge learning and old knowledge retention. This framework aims to continually accumulate knowledge and alleviate catastrophic forgetting guided by the Bayesian rule.
@@
4.2. Bayesian Guided Knowledge Evolution Learning
@@
where $\hat{\boldsymbol{\theta}}_{i,t}$ is the current embedding, and $\beta$ is the regularization hyperparameters. This term is weighted by the precision $\boldsymbol{\lambda}_{t-1}$, limiting changes to highly certain knowledge, thereby preserving old knowledge.
@@ 5.1. Experimental Setup
Baselines. To fully demonstrate the effectiveness of BAKE, we compare it with all recent CKGE models, including Fin-tune (Cui et al., 2023) (based on fine-tuning with new triples), IncDE (Liu et al., 2024a) and FastKGE (Liu et al., 2024b) (based on incremental distillation and LoRA fine-tuning), LKGE (Cui et al., 2023), EWC (Kirkpatrick et al., 2017), SI (Zenke et al., 2017), DiCGRL (Kou et al., 2020), EMR (Wang et al., 2019), CWR (Lopez-Paz and Ranzato, 2017), PNN (Rusu et al., 2016), FMR (Zhu et al., 2024), as well as energy-based and fine-grained token-driven models ETT-CKGE (Zhu et al., 2025), SAGE(Li et al., 2025b) and CLKGE (Cao, 2024).
@@ 5.2. Results and Analysis
Main Results and Analysis.
Experimental results are shown in the Table 1 and Table 2. Overall, our proposed BAKE method achieves state-of-the-art performance across all benchmark datasets and evaluation metrics (MRR, Hits@1/3/10), validating the effectiveness of the Bayesian-guided continual learning framework for the continual Knowledge Graph Embedding (CKGE) task. For large-scale initial knowledge graphs (e.g., FB-CKGE) and sparse or difficult scenarios (e.g., WN-CKGE), BAKE achieves state-of-the-art MRR and consistently maintains strong Hits, demonstrating robust knowledge acquisition while resisting catastrophic forgetting. These advantages are consistent with the design goal of combining sequential Bayesian updates with representation-level constraints.
@@ A.1. Detailed Experiment Settings
SAGE (Scale-Aware Gradual Evolution(Li et al., 2025b)). Adapts embedding size to update scale and uses dynamic distillation to balance retention and absorption, yielding robust gains across benchmarks.
CLKGE  (Cao, 2024). Learns an energy manifold for dynamic KGs and aligns old/new energies to mitigate forgetting while transferring knowledge forward.
@@
A.2. Related Work on Continual Learning
Continual learning (CL) enables models to learn from a continuous data stream, addressing the ”catastrophic forgetting” of old knowledge, a problem first identified in neural networks by McCloskey and Cohen(McCloskey and Cohen, 1989). The goal is to balance stability (retaining knowledge) with plasticity.

### v3

@@ Abstract.
Knowledge Graph; Knowledge Graph Embedding; Clustering; Continual Learning
copyright: acmlicenseddoi: 3774904.3792692journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792692isbn: 979-8-4007-2307-0/2026/04ccs: Computing methodologies Reasoning about belief and knowledge
@@ 1. Introduction
Knowledge graphs (KGs)(Liang et al., 2024b)(Ji et al., 2021) have become indispensable tools for representing structured information, supporting a wide range of knowledge‑driven applications such as LLM‑based question answering(Pan et al., 2024)(Xu et al., 2025), Blockchain(Song et al., 2025, 2022), and recommender systems(Zhao et al., 2024). Knowledge graph embedding (KGE) aims to encode entities and relations in a KG as continuous vectors, and the quality of these embeddings directly affects downstream performance. Classic KGE models (Bordes et al., 2013)(Sun et al., 2019)(Li et al., 2025a) perform excellently on static KGs, yet real‑world graphs usually evolve continuously. For example, the YAGO knowledge base evolved from YAGO3(Mahdisoltani et al., 2013) to YAGO4.5(Suchanek et al., 2024), with the number of entities soaring from about 4.6 million to 49 million and facts (triples) from 40 million to 132 million. For such continually evolving knowledge graphs (CKGs), existing KGE models must retrain the entire graph after each update—a computationally prohibitive and impractical strategy for large‑scale graphs like YAGO. Hence, this work focuses on CKGE models capable of learning CKGs efficiently.
@@ 2.3. Continual learning and Bayesian neural networks
Exact inference is usually intractable for deep networks, so an independent Gaussian (mean‑field) approximation is commonly adopted (Lee et al., 2024b; Ahn et al., 2019): $q_{t}(\boldsymbol{\omega})=\prod_{i=1}^{P}\mathcal{N}\!\bigl(\omega_{i};\mu_{t,i},\sigma_{t,i}^{2}\bigr)$, with $\theta_{t}=(\boldsymbol{\mu}_{t},\boldsymbol{\sigma}_{t})$ taken as the learnable variational parameters. The independence assumption greatly simplifies the KL‐divergence computation and the reparameterization gradient estimator, enabling iterative posterior updates. In practice, the reparameterization trick $\omega_{i}=\mu_{t,i}+\sigma_{t,i}\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,1)$ is often employed to estimate gradients.
@@ 3.1. Framework Overview
The overall framework of BAKE is shown in Figure 2. Its core idea is to formalize the continual knowledge graph embedding (CKGE) problem as a serialized Bayesian inference process. Given a knowledge graph sequence $\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{T}$ that evolves over time, the model receives a set of newly added triples $\Delta\mathcal{T}_{t}$ at each time snapshot $t$. BAKE works together through the following three modules: First, Bayesian-guided knowledge evolution uses the parameter posterior distribution of the previous snapshot t1t-1 as a prior, integrates new knowledge and quantifies uncertainty through Bayesian updating; second, the continual clustering module introduces fairness regularization to maintain the semantic consistency of the embedding space; finally, the joint optimization objective integrates the above modules to achieve a balance between new knowledge learning and old knowledge retention. This framework aims to continually accumulate knowledge and alleviate catastrophic forgetting guided by the Bayesian rule.
@@
3.2. Bayesian‐Guided Knowledge Evolution Learning
@@
where $\hat{\boldsymbol{\theta}}_{i,t}$ is the current embedding, and $\beta$ is the regularization hyperparameter. This term is weighted by the precision $\boldsymbol{\lambda}_{t-1}$, limiting changes to highly certain knowledge, thereby preserving old knowledge.
@@ 4.1. Experimental Setup
Baselines. To fully demonstrate the effectiveness of BAKE, we compare it with all recent CKGE models, including Fine-tune (Cui et al., 2023) (based on fine-tuning with new triples), IncDE (Liu et al., 2024a) and FastKGE (Liu et al., 2024b) (based on incremental distillation and LoRA fine-tuning), LKGE (Cui et al., 2023), EWC (Kirkpatrick et al., 2017), SI (Zenke et al., 2017), DiCGRL (Kou et al., 2020), EMR (Wang et al., 2019), CWR (Lopez-Paz and Ranzato, 2017), PNN (Rusu et al., 2016), FMR (Zhu et al., 2024), as well as energy-based and fine-grained token-driven models ETT-CKGE (Zhu et al., 2025), SAGE(Li et al., 2025c) and CLKGE (Cao, 2024).
@@ 4.2. Results and Analysis
Main Results and Analysis. Experimental results are shown in the Table 1 and Table 2. Overall, our proposed BAKE method achieves state-of-the-art performance across all benchmark datasets and evaluation metrics (MRR, Hits@1/3/10), validating the effectiveness of the Bayesian-guided continual learning framework for the continual Knowledge Graph Embedding (CKGE) task. For large-scale initial knowledge graphs (e.g., FB-CKGE) and sparse or difficult scenarios (e.g., WN-CKGE), BAKE achieves state-of-the-art MRR and consistently maintains strong Hits, demonstrating robust knowledge acquisition while resisting catastrophic forgetting. These advantages are consistent with the design goal of combining sequential Bayesian updates with representation-level constraints.
@@ 5. Conclusion
Acknowledgements.
This work was supported by the Key Program of the National Natural Science Foundation of China (Grant No. 62436006).
@@ A.1. Detailed Experiment Settings
SAGE (Scale-Aware Gradual Evolution(Li et al., 2025c)). Adapts embedding size to update scale and uses dynamic distillation to balance retention and absorption, yielding robust gains across benchmarks.
CLKGE (Continual Learning Knowledge Graph Embeddings (Cao, 2024)). Learns an energy manifold for dynamic KGs and aligns old/new energies to mitigate forgetting while transferring knowledge forward.
@@
A.1.4. MRR Metric.
Mean Reciprocal Rank (MRR) evaluates the ranking quality by considering the reciprocal of the rank position of the correct answer. Formally, given a query set $\mathcal{Q}$ at snapshot $i$, let $\mathrm{rank}_{q}$ denote the rank of the ground-truth entity (or answer) for query $q\in\mathcal{Q}$. The MRR score is defined as:
(11) $\mathrm{MRR}=\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\frac{1}{\mathrm{rank}_{q}}.$
Higher MRR indicates correct answers rank nearer the top.
A.1.5. Hits@N Metric.
Hits@N (also denoted as Hits@K) measures the proportion of queries whose correct answers appear in the top-$N$ ranked results. Specifically, for each query $q\in\mathcal{Q}$, we define an indicator function $\mathbb{I}(\mathrm{rank}_{q}\leq N)$, then Hits@N is computed as:
(12) $\mathrm{Hits@}N=\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\mathbb{I}(\mathrm{rank}_{q}\leq N).$
A larger Hits@N indicates that the model retrieves correct answers within the top-$N$ positions more frequently.
@@
A.2. Related Work
A.2.1. Knowledge Graph Embedding
Early KGE models mainly follow the translation paradigm, such as TransE(Bordes et al., 2013) and TransH(Wang et al., 2014) in the real vector space, and their extensions to complex/quaternion spaces, including RotatE(Sun et al., 2019), QuatE(Zhang et al., 2019), HAKE(Zhang et al., 2020), SATKGC(Ko et al., 2025), MRME(Li et al., 2025b), SymAgent(Liu et al., 2025), HySAE(Li et al., 2025d), and GIE(Cao et al., 2022). These approaches introduced the idea of representing a triple $(h,r,t)$ through simple algebraic operations in a vector space. Since then, both the efficiency and the expressive power of KGE models have improved along multiple dimensions. RecPiece(Liang et al., 2024a) clusters relation centres to select anchors, reducing memory consumption while improving link prediction accuracy; RAA‑KGC(Yuan et al., 2025a) pulls queries of a pretrained language model toward relation‑aware neighbour anchors to construct more discriminative embeddings; FedE(Zhou et al., 2024) pioneers a poisoning attack on federated KG embedding, using inferred relations to craft stealthy updates that induce specific errors without degrading main task performance. Recently, KGE research has begun to leverage diffusion models to recast the task as conditional entity generation. For example, FDM(Long et al., 2024) trains a diffusion model that outputs the distribution of tail entities conditioned on the head entity and relation, thereby capturing multimodal relational semantics without relying on complicated probabilistic mechanisms. S2DN(Ma et al., 2025) proposes a semantic-aware denoising network and filters out unreliable interactions to enhance the structural reliability of target links.
A.2.2. Continual Knowledge Graph Embedding
As knowledge graphs (KGs) evolve, embedding models must learn new facts while avoiding catastrophic forgetting of previously acquired knowledge. Research on continual knowledge graph embedding (CKGE)(Li et al., 2024) has progressed notably in recent years. IncDE (Liu et al., 2024a) orders new triples hierarchically and employs layer-wise incremental distillation to preserve old knowledge; FastKGE (Liu et al., 2024b) accelerates updates with an adaptive low-rank adapter (IncLoRA), substantially shortening training time while isolating parameters for new snapshots; LKGE (Cui et al., 2023) combines a masked autoencoder with transfer regularization to extend embeddings across multiple snapshots; ETT-CKGE (Zhu et al., 2025) replaces costly node scoring with lightweight task-driven labels and aligns KG snapshots through simple matrix operations; CLKGE (Cao, 2024) couples continual learning with an energy-based manifold so that new and old knowledge reinforce each other; CMKGE (Song et al., 2024) introduces a biologically inspired dual inhibition–excitation mask to balance plasticity and stability; FMR (Zhu et al., 2024) rotates the parameter space based on Fisher information and assigns flexible regularization to further mitigate forgetting; ERPP (Yang et al., 2025) transfers relational paths across snapshots to shift from catastrophic forgetting to knowledge accumulation; CFKGC (Li et al., 2024) introduces model-level modulation to mitigate catastrophic forgetting while leveraging multi-view relation augmentation to better learn scarce emerging relations; and SAGE (Li et al., 2025c) adaptively expands embedding dimensions and uses dynamic distillation for CKGE.
A.2.3. Continual Learning
Continual learning (CL)(Wang et al., 2024) enables models to learn from a continuous data stream, addressing the ”catastrophic forgetting” of old knowledge, a problem first identified in neural networks by McCloskey and Cohen(McCloskey and Cohen, 1989). Unlike time series prediction (Zhang et al., 2025), the goal is to balance stability (retaining knowledge) with plasticity.
@@ A.3. BAKE model training process pseudocode
A.4. Additional Discussion
A.4.1. End-user tasks and application relevance.
Knowledge graph embeddings are a foundational representation for many knowledge-driven applications, including LLM-based question answering and recommender systems, where embedding quality directly affects downstream performance (Yuan et al., 2025b). In continual settings, these applications require models to rapidly absorb newly emerging facts while maintaining previously learned semantics. In this work, we follow the standard CKGE protocol by evaluating models on continual link prediction, i.e., ranking candidate entities for incomplete queries such as $(h,r,?)$ while learning new knowledge without forgetting past knowledge. This evaluation is not only widely adopted in prior CKGE studies but also closely aligned with real-world needs because link prediction is the core operation behind KG completion and retrieval.
A.4.2. Runtime and computational efficiency.
We acknowledge the importance of computational efficiency in continual learning and complement our accuracy evaluation with an efficiency discussion. BAKE introduces lightweight overhead on top of the backbone KGE training: the sequential Bayesian posterior-as-prior update is implemented in closed form with per-dimension operations, as shown by the element-wise update rules and the element-wise product operator. The continual clustering component maintains centroids via a momentum update, which is performed once per epoch in our training procedure. In our implementation, BAKE’s training time per snapshot is within about $1.2\times$ of IncDE; the main extra cost comes from centroid maintenance and reassignment, while Bayesian updates are negligible compared with standard mini-batch KGE optimization. All experiments are conducted under the same hardware setting (8 NVIDIA Tesla V100 GPUs).
A.4.3. Choice of TransE as the base scorer.
We use TransE as the base scorer to generate snapshot-specific embedding observations, following the standard KGE training objective with $f(h,r,t)=\|h+r-t\|_{2}$ (Bordes et al., 2013). This design ensures direct comparability to many recent CKGE baselines that adopt the same backbone, while allowing us to isolate the contribution of the proposed continual Bayesian framework. Importantly, BAKE does not rely on TransE-specific assumptions: in our formulation, the embeddings $\hat{\Theta}_{t}$ obtained from the snapshot learner are treated as observations used for Bayesian updating, and the posterior update itself is model-agnostic as long as a snapshot learner can produce $\hat{\Theta}_{t}$. Therefore, BAKE can, in principle, be combined with more expressive KGE scorers (e.g., RotatE, QuatE)(Sun et al., 2019; Zhang et al., 2019), while we intentionally keep the base scorer fixed in this paper to provide a controlled and fair comparison across continual-learning strategies.